OpenStudy (anonymous):

Can someone tell me if I am doing this the correct way? I have to figure out if it is always, sometimes, or never true. 2(8x-4)-2x (is less than or equal to) 14x+12 I know you distribute making it: 16x-8-2x (is less than or equal to) 14x+12 then, I got 14x-8 (is less than or equal to) 14x+12 What is the next step? Do I subtract 14x from both sides and get -8 is less than or equal to 12?

OpenStudy (texaschic101):

2(8x - 4)- 2x <= 14x + 12 16x - 8 - 2x <= 14x + 12 14x - 8 <= 14x + 12 (subtract 14x from both sides) 14x - 14x - 8 <= 12 (now add 8 to both sides) 0 - 8 + 8 <= 12 + 8 0 <= 20 test it... 14x - 8 <= 14x + 12 (test number 5) 14(5) - 8 <= 14(5) + 12 70 - 8 <= 70 + 2 62 <= 72 (true) use -5 as your test number 14x - 8 <= 14x + 12 14(-5) - 8 <= 14(-5) + 12 -70 - 8 <= -70 + 12 -78 <= - 58 (true) It is ALWAYS TRUE